2(n+5)=16 what is the valu of n?

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

2(n+5)=16
/2
(n+5)=8
-5
n+5-5 = 8-5
n=3

Answer by stanbon(75887)

2(n+5)=16 what is the value of n?
-----
n+5 = 8
----
n = 3
=============
Cheers,
Stan H.